A box contains $6.25 in nickels, dimes, and quarters. There are 85 coins in all, with three times as many nickels as dimes. How many quarters are there.
Answer by Greenfinch(383)

no of dimes is X, no of nickels is 3X, no of quarters is 85 - 4X
so 10X + (3Xx5) + (85 - 4x)x25 = 625 (do it all in cents)
10X + 15X + 2125 -100X = 625
-75X = -1500
X = 20,
so 20 dimes, 60 nickels and 5 quarters
To check 20 x 10 = 200
60 x 5 = 300
5 x 25 = 125
total = 625 or $6.25